OpenStudy (anonymous):

3. Write an equation of a line in point slope form that has a slope of -3 and passes through the point 3,-4

OpenStudy (anonymous):

y-y1=m(x-x1) m=-3 x1=3 y1=-4 Plug the numbers in

OpenStudy (anonymous):